# Existence for Stable Rotating Star-Planet Systems

### v1

@@ 4 Existence of Constrained Minimizers of $E_{J}(\rho)$ on ${W}_{m}$
We want to prove the existence of “double constrained” minimizers via direct method [17, Theorem 1.15]. Here, “double constrained” means $\rho(m)\in W_{m}$ and $\rho(m)\leq R$ a.e.. To do that we first show the lower semicontinuity of $E_{J}$ by reformulating arguments in [4] where the authors only consider non-rotating one body case.
With an abuse of notation, we denote sequences $\left\{\rho_{n,R}\right\}\subset W_{m,R}$ and their limit $\rho_{R}$. Note their total mass $\int\rho_{n,R}$} is always one.
@@
If $\left\{\rho_{n,R}\right\}\subset W_{m,R}$ and $\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}$, then for almost every $x\in\mathbb{R}^{3},0\leq\rho_{n,R}\leq$ $R$ and up to subsequence $\rho_{n,R}\rightarrow\rho_{R}$. Then $0\leq\rho_{R}\leq R$ almost everywhere. Since $\Omega:=$ $\Omega_{m}\cup\Omega_{1-m}$ is bounded, by Hölder’s inequality we know $\left\|\rho_{n,R}-\rho_{R}\right\|_{L^{1}}\leq C\|\rho_{n,R}$- $\rho_{R}\|_{L^{\frac{4}{3}}}\rightarrow 0$. It implies $\rho_{R}\in W_{m,R}$ thus $W_{m,R}$ is closed. Easy to check it is also a convex subset of $L^{\frac{4}{3}}$. Therefore, $W_{m,R}$ is weakly closed.
@@
To proof the lower semicontinuity, we recall $E_{J}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}+T_{J}(\rho)=\int A(\rho)$- $\frac{\int V_{\rho}\cdot\rho}{2}+T_{J}(\rho)=\int\frac{K(\rho(x))^{\gamma}}{\gamma-1}d^{3}x-\frac{1}{2}\iint\frac{\rho(x)\cdot\rho(y)}{|x-y|}d^{3}xd^{3}y+\frac{J^{2}}{2(I(\rho))}$. We first notice $A(\rho)$ is a convex function, so $\int A(\rho)$ is a convex functional. Then for each constant $k,\left\{\rho_{R}\in\right$. $\left.W_{m,R}\mid\int A\left(\rho_{R}\right)\leq k\right\}$ is convex. $\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}$ implies almost everywhere convergence up to a subsequence, and by construction we know $\rho_{n,R}\leq R\cdot 1_{\Omega}$, here $1_{\Omega}$ is the indicator function of compact set $\Omega$. By (2) we know $A(0)=0$, $A(\cdot)$ is non-decreasing, and thus $A(\rho_{n,R})\leq A(R)\cdot 1_{\Omega}\in L^{1}$. Therefore, if $\int A\left(\rho_{n,R}\right)\leq k$ then $\int A\left(\rho_{R}\right)\leq k$ by dominated convergence theorem or Fatou’s lemma, which means $\left\{\rho_{R}\in W_{m,R}\mid\int A\left(\rho_{R}\right)\leq k\right\}$ is not only convex but also closed. Thus, it is also weakly closed and then $\int A(\rho)$ is a weakly lower semicontinuous functional. Next, thanks to Proposition A.5, we know $\mathcal{V}:\rho\rightarrow V_{\rho}$ is compact from $L^{\frac{4}{3}}(\Omega)$ to $L^{r}(\Omega)$ for any $1\leq r<12$, thus $\mathcal{V}$ is completely continuous. Take $r=11$, then $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}(\Omega)$ implies $V_{\rho_{n,R}}\rightarrow V_{\rho_{R}}$ in $L^{11}(\Omega)$. Since $\Omega$ is bounded, by Holder’s inequality we know $V_{\rho_{n,R}}\rightarrow V_{\rho_{R}}$ in $L^{4}(\Omega)$. Since $\rho_{R}$ and $\rho_{n,R}$ vanish outside $\Omega$, we have
$\left|G\left(\rho_{R},\rho_{R}\right)-G\left(\rho_{n,R},\rho_{n,R}\right)\right|=$ $\left|\int_{\mathbb{R}^{3}}V_{\rho_{R}}\cdot\rho_{R}-\int_{\mathbb{R}^{3}}V_{\rho_{n,R}}\cdot\rho_{n,R}\right|$
$=$ $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\right|+\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}-\int_{\Omega}V_{\rho_{n,R}}\cdot\rho_{n,R}\right|$
$\leq$ $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\right|+\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|V_{\rho_{R}}-V_{\rho_{n,R}}\right\|_{L^{4}(\Omega)}$
@@
The last inequality comes from the Hölder’s inequality. Note $V_{\rho_{R}}\in L^{11}(\Omega)\subset L^{4}(\Omega)=\left(L^{\frac{4}{3}}(\Omega)\right)^{*}$, together with $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ we know $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\right|\rightarrow 0$. Note $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ also implies $\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq C$, thus $\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|V_{\rho_{R}}-V_{\rho_{n,R}}\right\|_{L^{4}(\Omega)}\rightarrow 0$ due to $V_{\rho_{n,R}}\rightarrow V_{\rho_{R}}$ in $L^{4}(\Omega)$. Hence we have $G\left(\rho_{n,R},\rho_{n,R}\right)\rightarrow G\left(\rho_{R},\rho_{R}\right)$. Finally, if $\rho\in W_{m,R}$, then $\bar{x}(\rho)=\frac{\int_{\mathbb{R}^{3}}\rho(x)x}{\int_{\mathbb{R}^{3}}\rho}=\int_{\Omega}\rho(x)x$, and $\rho_{n,R}\rightharpoonup\rho_{R}$ implies $\bar{x}\left(\rho_{n,R}\right)\rightarrow\bar{x}\left(\rho_{R}\right)$ since $f(x)=x\in L^{\infty}\left(\Omega;\mathbb{R}^{3}\right)\subset\left(L^{\frac{4}{3}}\left(\Omega;\mathbb{R}^{3}\right)\right)^{*}$. Moreover,
@@
$=\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-\int_{\Omega}\rho_{R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right|$
$\leq\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right|+\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)-\int_{\Omega}\rho_{R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right|$
$\leq\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right\|_{L^{4}(\Omega)}+\left|\int_{\Omega}\left(\rho_{n,R}(x)-\rho_{R}(x)\right)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right|$
By an argument similar to the above, we know $r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)$ and $r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)$ are in $L^{\infty}(\Omega)\subset$ $L^{4}(\Omega)=\left(L^{\frac{4}{3}}(\Omega)\right)^{*}$. $\bar{x}\left(\rho_{n,R}\right)\rightarrow\bar{x}\left(\rho_{R}\right)$ means $r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)\rightarrow r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)$ everywhere, thus by dominated convergence theorem we know $\|r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\|_{L^{4}(\Omega)}\rightarrow 0$. $\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ implies weak convergence in $L^{\frac{4}{3}}(\Omega)$, $\left\|\rho_{n,R}\right\|_{L^{3}(\Omega)}\leq C$, then we have $I\left(\rho_{n,R}\right)\rightarrow I\left(\rho_{R}\right),\frac{J^{2}}{2\left(I\left(\rho_{n,R}\right)\right)}\rightarrow\frac{J^{2}}{2\left(I\left(\rho_{R}\right)\right)}$. Collect the results above we obtain $E_{J}\left(\rho_{R}\right)\leq\liminf\limits_{n\rightarrow\infty}E_{J}\left(\rho_{n,R}\right)$. ∎
@@
Fix $J>0$, give $m>0,R>\frac{384}{\pi\eta^{3}}$, there is a $\rho_{R}\in W_{m,R}$ which minimizes $E_{J}(\rho)$ on $W_{m,R}$. Any such $\rho_{R}$ is continuous on $\Omega=\Omega_{m}\cup\Omega_{1-m}$. Moreover, there are some constants $\lambda_{m,R}$ and $\lambda_{1-m,R}$ such that
@@
Given a minimizing sequence $\left\{\rho_{n,R}\right\}$ with $\lim\limits_{n\rightarrow\infty}E\left(\rho_{n,R}\right)=\inf\limits_{\rho\in W_{m,R}}E(\rho)$. By Interpolation Inequality Proposition A.2 we know $W_{m,R}$ is bounded in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$. Since $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ is reflexive, $\exists\rho_{R}\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ such that $\rho_{n,R}\rightharpoonup\rho_{R}$ up to subsequence [8, Theorem 3.18]. By Lemma 4.3 $W_{m,R}$ is weakly compact in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right),\rho_{R}\in W_{m,R}$, and $\inf\limits_{\rho\in W_{m,R}}E(\rho)\leq E\left(\rho_{R}\right)\leq\lim\limits_{n\rightarrow\infty}E\left(\rho_{n,R}\right)=\inf\limits_{\rho\in W_{m,R}}E(\rho)$, which means $\rho_{R}$ is a minimizer. A slight modification of the proof of [4, Lemma 2], [15, Proposition 5.8] and [16, Lemma 2.21], which means we need to consider the perturbations have compact supports in $\Omega$, shows that (33) and (34) hold for almost all points which are mentioned in (33) (34) with suitable $\lambda_{m,R}$ or $\lambda_{1-m,R}$. The continuity of $\rho_{R}$ on $\Omega$ then follows as in the proof of Theorem 3.6 (see e.g. [15, Theorem 2.25] and [16, Remark 2.15]), where the idea is making use of (33) (34) (35) and improving the regularity of potential $V_{\rho_{R}}$ and density $\rho_{R}$ in turns (bootstrap). Then we know (33) and (34) hold indeed for all points we considered. ∎
@@
Another type of bootstrap, commonly used in fluid dynamics and general relativity, refers to bootstrapping with respect to time. This differs from the regularity bootstrap discussed in our paper.
@@
We have $\forall R\in\left[R_{0},\infty\right),\rho_{R_{0}}\in W_{m,R}$, then $E\left(\rho_{R}\right)\leq E\left(\rho_{R_{0}}\right)<\infty$. Then, one can adapt the arguments from the non-rotating case—see, for instance, [16, Lemma 2.21] or [4, Section 6]—to show that the $L^{\frac{4}{3}}$ norm of $\rho_{R}$ is actually bounded by a constant depends only on $E\left(\rho_{R_{0}}\right)$. ∎
Deducing the uniform boundedness in $L^{\infty}$ from that in $L^{\frac{4}{3}}$ can be accomplished in several ways. For instance, one may refer to the methods of Li in [38, Proposition 1.4], as McCann mentioned in [46, Section 6]. Li’s method [38] can be applied to obtain uniform boundedness in $L^{\infty}$ for $\gamma>\frac{4}{3}$ by choosing an appropriate test function, although it does not first verify whether Lagrange multiplier is negative — Li verified this after the $L^{\infty}$ bound was obtained in [38].
Here, we present an alternative approach, which is inspired by the method in [4]. Both methods rely on the result that velocity $v^{2}(x)=\frac{J^{2}r^{2}(x-\bar{x}(\rho))}{I^{2}(\rho)}$ is sufficiently small. Under assumption $\gamma>\frac{3}{2}$, we first prove that Lagrange multipliers are negative, using the idea from [4, Lemma 5] and [38, Proposition 2.5].

### v2

@@ 4 Existence of Constrained Minimizers of $E_{J}(\rho)$ on ${W}_{m}$
We want to prove the existence of “double constrained” minimizers via direct method [12, Theorem 1.15]. Here, “double constrained” means $\rho(m)\in W_{m}$ and $\rho(m)\leq R$ a.e.. To do that, we first show the lower semicontinuity of $E_{J}$ by reformulating arguments in [4] where the authors only consider non-rotating one-body case.
With an abuse of notation, we denote sequences in $W_{m,R}$ by $\left\{\rho_{n,R}\right\}$ and their limit by $\rho_{R}$. Note their total mass ${\int_{\mathbb{R}^{3}}\rho_{n,R}}\,dx$ is always one.
@@
If $\left\{\rho_{n,R}\right\}\subset W_{m,R}$ and $\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}$, then for almost every $x\in\mathbb{R}^{3},0\leq\rho_{n,R}\leq$ $R$ and up to a subsequence $\rho_{n,R}\rightarrow\rho_{R}$. Then $0\leq\rho_{R}\leq R$ almost everywhere. Since $\Omega:=$ $\Omega_{m}\cup\Omega_{1-m}$ is bounded, by Hölder’s inequality we know $\left\|\rho_{n,R}-\rho_{R}\right\|_{L^{1}}\leq C\|\rho_{n,R}$- $\rho_{R}\|_{L^{\frac{4}{3}}}\rightarrow 0$. It implies $\rho_{R}\in W_{m,R}$ thus $W_{m,R}$ is closed. Easy to check it is also a convex subset of $L^{\frac{4}{3}}$. Therefore, $W_{m,R}$ is weakly closed.
To prove the lower semicontinuity, we recall
$E_{J}(\rho)$ $=U(\rho)-\frac{G(\rho,\rho)}{2}+T_{J}(\rho)$
$=\int_{\mathbb{R}^{3}}A(\rho)\,dx-\frac{\int_{\mathbb{R}^{3}}V_{\rho}\cdot\rho\,dx}{2}+T_{J}(\rho)$
$=\int_{\mathbb{R}^{3}}\frac{K(\rho(x))^{\gamma}}{\gamma-1}\,dx-\frac{1}{2}\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\rho(x)\cdot\rho(y)}{|x-y|}\,dx\,dy+\frac{J^{2}}{2(I(\rho))}$
@@
We first notice $A(\rho)$ is a convex function, so $\int_{\mathbb{R}^{3}}A(\rho)\,dx$ is a convex functional. Then for each constant $k$, $\left\{\rho_{R}\in\right$. $\left.W_{m,R}\mid\int_{\mathbb{R}^{3}}A\left(\rho_{R}\right)\,dx\leq k\right\}$ is convex. $\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}$ implies almost everywhere convergence up to a subsequence, and by construction we know $\rho_{n,R}\leq R\cdot 1_{\Omega}$, here $1_{\Omega}$ is the indicator function of compact set $\Omega$. By (2) we know $A(0)=0$, $A(\cdot)$ is non-decreasing, and thus $A(\rho_{n,R})\leq A(R)\cdot 1_{\Omega}\in L^{1}$. Therefore, if $\int_{\mathbb{R}^{3}}A\left(\rho_{n,R}\right)\,dx\leq k$ then $\int_{\mathbb{R}^{3}}A\left(\rho_{R}\right)\,dx\leq k$ by dominated convergence theorem or Fatou’s lemma, which means $\left\{\rho_{R}\in W_{m,R}\mid\int_{\mathbb{R}^{3}}A\left(\rho_{R}\right)\,dx\leq k\right\}$ is not only convex but also closed. Thus, it is also weakly closed and then $\int_{\mathbb{R}^{3}}A(\rho)\,dx$ is a weakly lower semicontinuous functional.
Next, we know $\mathcal{V}:\rho\rightarrow V_{\rho}$ is compact from $L^{\frac{4}{3}}(\Omega)$ to $L^{r}(\Omega)$ for any $1\leq r<12$ (see, e.g., [11, Appendix A]), thus $\mathcal{V}$ is completely continuous. Take $r=11$, then $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}(\Omega)$ implies $V_{\rho_{n,R}}\rightarrow V_{\rho_{R}}$ in $L^{11}(\Omega)$. Since $\Omega$ is bounded, by Holder’s inequality we know $V_{\rho_{n,R}}\rightarrow V_{\rho_{R}}$ in $L^{4}(\Omega)$. Since $\rho_{R}$ and $\rho_{n,R}$ vanish outside $\Omega$, we have
$\left|G\left(\rho_{R},\rho_{R}\right)-G\left(\rho_{n,R},\rho_{n,R}\right)\right|=$ $\left|\int_{\mathbb{R}^{3}}V_{\rho_{R}}\cdot\rho_{R}\,dx-\int_{\mathbb{R}^{3}}V_{\rho_{n,R}}\cdot\rho_{n,R}\,dx\right|$
$=$ $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}\,dx-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\,dx\right|+\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\,dx-\int_{\Omega}V_{\rho_{n,R}}\cdot\rho_{n,R}\,dx\right|$
$\leq$ $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}\,dx-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\,dx\right|+\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|V_{\rho_{R}}-V_{\rho_{n,R}}\right\|_{L^{4}(\Omega)}$
@@
The last inequality comes from Hölder’s inequality. Note $V_{\rho_{R}}\in L^{11}(\Omega)\subset L^{4}(\Omega)=\left(L^{\frac{4}{3}}(\Omega)\right)^{*}$, together with $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ we know $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}\,dx-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\,dx\right|\rightarrow 0$. Note $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ also implies $\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq C$, thus $\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|V_{\rho_{R}}-V_{\rho_{n,R}}\right\|_{L^{4}(\Omega)}\rightarrow 0$ due to $V_{\rho_{n,R}}\rightarrow V_{\rho_{R}}$ in $L^{4}(\Omega)$. Hence we have $G\left(\rho_{n,R},\rho_{n,R}\right)\rightarrow G\left(\rho_{R},\rho_{R}\right)$.
Finally, if $\rho\in W_{m,R}$, then $\bar{x}(\rho)=\frac{\int_{\mathbb{R}^{3}}\rho(x)x\,dx}{\int_{\mathbb{R}^{3}}\rho\,dx}=\int_{\Omega}\rho(x)x\,dx$, and $\rho_{n,R}\rightharpoonup\rho_{R}$ implies $\bar{x}\left(\rho_{n,R}\right)\rightarrow\bar{x}\left(\rho_{R}\right)$ since $f(x)=x\in L^{\infty}\left(\Omega;\mathbb{R}^{3}\right)\subset\left(L^{\frac{4}{3}}\left(\Omega;\mathbb{R}^{3}\right)\right)^{*}$. Moreover,
@@
$=\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)\,dx-\int_{\Omega}\rho_{R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\,dx\right|$
$\leq\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)\,dx-\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\,dx\right|$
$+\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\,dx-\int_{\Omega}\rho_{R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\,dx\right|$
$\leq\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right\|_{L^{4}(\Omega)}+\left|\int_{\Omega}\left(\rho_{n,R}(x)-\rho_{R}(x)\right)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\,dx\right|$
By an argument similar to the above, we know $r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)$ and $r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)$ are in $L^{\infty}(\Omega)\subset$ $L^{4}(\Omega)=\left(L^{\frac{4}{3}}(\Omega)\right)^{*}$. $\bar{x}\left(\rho_{n,R}\right)\rightarrow\bar{x}\left(\rho_{R}\right)$ means $r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)\rightarrow r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)$ everywhere, thus by dominated convergence theorem we know $\|r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\|_{L^{4}(\Omega)}\rightarrow 0$. $\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ implies weak convergence in $L^{\frac{4}{3}}(\Omega)$, $\left\|\rho_{n,R}\right\|_{L^{3}(\Omega)}\leq C$, then we have $I\left(\rho_{n,R}\right)\rightarrow I\left(\rho_{R}\right)$, and then $\frac{J^{2}}{2\left(I\left(\rho_{n,R}\right)\right)}\rightarrow\frac{J^{2}}{2\left(I\left(\rho_{R}\right)\right)}$.
Collecting the results above, we obtain $E_{J}\left(\rho_{R}\right)\leq\liminf\limits_{n\rightarrow\infty}E_{J}\left(\rho_{n,R}\right)$. ∎
@@
Fix $J>0$, given $m>0,R>\frac{384}{\pi\eta^{3}}$, there is a $\rho_{R}\in W_{m,R}$ which minimizes $E_{J}(\rho)$ on $W_{m,R}$. Any such $\rho_{R}$ is continuous on $\Omega=\Omega_{m}\cup\Omega_{1-m}$. Moreover, there are some constants $\lambda_{m,R}$ and $\lambda_{1-m,R}$ such that
@@
Given a minimizing sequence $\left\{\rho_{n,R}\right\}$ with $\lim\limits_{n\rightarrow\infty}E\left(\rho_{n,R}\right)=\inf\limits_{\rho\in W_{m,R}}E(\rho)$. By Interpolation Inequality [6, Section 4.2] we know $W_{m,R}$ is bounded in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$. Since $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ is reflexive, $\exists\rho_{R}\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ such that $\rho_{n,R}\rightharpoonup\rho_{R}$ up to subsequence [6, Theorem 3.18]. By Lemma 4.3 $W_{m,R}$ is weakly compact in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right),\rho_{R}\in W_{m,R}$, and
$\inf\limits_{\rho\in W_{m,R}}E(\rho)\leq E\left(\rho_{R}\right)\leq\lim\limits_{n\rightarrow\infty}E\left(\rho_{n,R}\right)=\inf\limits_{\rho\in W_{m,R}}E(\rho)$
It means $\rho_{R}$ is a minimizer. A slight modification of the proof of [4, Lemma 2], [10, Section 5] and [11, Section 2], which means we need to consider the perturbations have compact supports in $\Omega$, shows that (34) and (35) hold for almost all points which are mentioned in (34) (35) with suitable $\lambda_{m,R}$ or $\lambda_{1-m,R}$. The continuity of $\rho_{R}$ on $\Omega$ then follows as in the proof of Theorem 3.6 (see, e.g., subsection 5.2 of this paper, [4], [10, Section 2] and [11, Section 2]), where the idea is making use of (34) (35) (36) and improving the regularity of potential $V_{\rho_{R}}$ and density $\rho_{R}$ in turns (bootstrap method). Then we know (34) and (35) hold indeed for all points we considered. ∎
@@
Since we consider potential and density alternatively, we call such method the bootstrap method. However, another type of bootstrap method, commonly used in fluid dynamics and general relativity, refers to bootstrapping with respect to time. It differs from the regularity bootstrap discussed in this paper.
@@
We have $\forall R\in\left[R_{0},\infty\right),\rho_{R_{0}}\in W_{m,R}$, then $E\left(\rho_{R}\right)\leq E\left(\rho_{R_{0}}\right)<\infty$. Then, one can adapt the arguments from the non-rotating case—see, for instance, [11, Section 2] or [4, Section 6]—to show that the $L^{\frac{4}{3}}$ norm of $\rho_{R}$ is actually bounded by a constant that depends only on $E\left(\rho_{R_{0}}\right)$. ∎
Deducing the uniform boundedness in $L^{\infty}$ from that in $L^{\frac{4}{3}}$ can be accomplished in several ways. For instance, one may refer to the methods of Li in [18, Proposition 1.4], as McCann mentioned in [24, Section 6]. Li’s method [18] can be applied to obtain uniform boundedness in $L^{\infty}$ for $\gamma>\frac{4}{3}$ by choosing an appropriate test function, although it is not first verified that Lagrange multiplier is negative — Li verified this after the $L^{\infty}$ bound was obtained in [18].
Here, we present an alternative approach, which is inspired by the method in [4]. Both methods rely on the result that velocity $v^{2}(x)=\frac{J^{2}r^{2}(x-\bar{x}(\rho))}{I^{2}(\rho)}$ is sufficiently small. Under assumption $\gamma>\frac{3}{2}$, we first prove that Lagrange multipliers are negative, using the idea from [4, Lemma 5] and [18, Proposition 2.5].